how do i work out a word problem in algebra format. john is 4 years younger than his sister mary. the sum of their ages is 20. using algebra by letting marys age = x and constructing an equation. find johns age. i know the answer is 8. but i do not know how to write it down in algebra
Answer by Edwin McCravy(20086)

john is 4 years younger than his sister mary. the sum of their ages is 20.
Just rewrite those two sentences using the words 

1. John's age
2. Mary's age
3. equals
4. minus
5. plus

"John is 4 years younger than his sister Mary"

can be rewritten as

"John's age equals Mary's age minus 4 years"

then you can put in = and - and get:

John's age = Mary's age - 4

Now rewrite:

"The sum of their ages is 20" using the above 5 words:

as

"John's age plus Mary's age equals 20"

then you can put in + and = and get:

John's age + Mary's age = 20.

Now let's look at the two sentences:


John's age = Mary's age - 4
John's age + Mary's age = 20.

Now let Mary's age = M, then we have:

John's age = M - 4
John's age + M = 20.

Now let's rewrite the second sentence but instead of 
"John's age" let's substitute M - 4 since the first sentence
tells us that's what John's age equals to:


John's age + M = 20

becomes

M - 4 + M = 20

Now collect the two like terms M and M and get 2M

2M - 4 = 20

Now add 4 to both sides and get

    2M = 24

Now divide both sides by 2 and get

     M = 12

So Mary's age is 12.

Now since 

John's age = M - 4

John's age = 12 - 4

John's age = 8

So Mary is 12 and John is 8.

Check: Mary is 4 years older than John because 12 is 4 more than 8

The sum of their ages is 20 because the sum of 12 and 8 is 20.

So it's correct.  Study the above carefully and maybe you'll
start to see how to translate English into algebra. 

Edwin